Joshua Industries is considering a new project with cash inflows of $500,000 for the indefinite future. Cash costs are 60 percent of the cash inflows. The initial cost of the investment is $700,000. The tax rate is 15 percent and the unlevered cost of equity is 17 percent. The firm is financing $170,000 of the project cost with debt. What is the adjusted present value of the project? Question 26 options: $102,429.67 $98,311.16 $93,940.85 $106,470.59 $325,500.00
Answer #1
Base case NPV = -700000+(500000*40%*85%)/17% = $         3,00,000
PV interest tax shield = 170000*15% = $            25,500
APV $   3,25,500.00